# The Open Journal of Astrophysics

### v1

@@ 4 Discussion & Outlook
Our work goes beyond the existing literature by remeasuring the shapes and variables used in the morphology definitions in a consistent way between simulations. Although we aim to make a fair comparison between the different simulations, due to the variations in the distributions of the morphologically defining variables, as described in Section 3.1, finding a consistent way to define and compare the samples across simulations remains a challenge. The origin of the differences between the distributions of the variables and their redshift evolution is unclear, as underlying stellar mass distributions of the samples, boxsize of the simulation, resolution, sub-grid physics models and galaxy and halo finder algorithms could all be playing a role.
@@
The $\mathrm{BTR}$ (Appendix D) and $|v/\sigma|$ splits in this work are similar to the work of Tenneti et al. (2016), where both variables are used as a morphological split to measure the ellipticity-direction correlations of disks around ellipticals in MassiveBlack-II (Khandai et al., 2015) and Illustris (Genel et al., 2014; Vogelsberger et al., 2014). These measurements are compared to those in Chisari et al. (2015), where $|v/\sigma|$ is used as a morphological proxy to measure the IA correlations of disks around ellipticals in Horizon-AGN. Tenneti et al. (2016) measure a positive (radial) disk around elliptical correlation in MassiveBlack-II and Illustris for selections baded on both $\mathrm{BTR}$ and $|v/\sigma|$ and show that while the amplitude changes with disk definition, the sign is always positive.
@@
Although this work gives insight into under what circumstances a negative disk signal is measured, further sub-selections of the galaxy samples and larger statistics could help unveil this. Measuring the same signal in larger boxes, such as the Horizon Run 5 (Lee et al., 2021), COLIBRE and MilleniumTNG , would also be very useful to continue this research. The effects of boxsize and resolution could also be explored further within these projects, as there will be multiple boxsizes and resolutions available that meet the minimum requirements to measure statistically significant correlations.
Another area to expand this research to would be the observational realm. Currently, little is known about how alignments respond to the kinematic properties of galaxies, as the measurement of $|v/\sigma|$ are very difficult to obtain observationally and only available for small samples (Pulsoni et al., 2020). Expanding works like Pulsoni et al. (2020) that compare kinematic properties of galaxies between hydrodynamical simulations and observations for small samples, will provide valuable insight on the accuracy of the models we currently employ.
This work highlights the importance of consistency when comparing between different disk and elliptical alignments under varying circumstances. Although hydrodynamical simulations agree on many features of the alignment signal, details like sample selection choices or measurement methodologies can make a substantive difference in outcome. Furthermore, this works calls for caution in weak lensing studies that restrict to blue galaxies to mitigate IA effects (e.g. McCullough et al., 2024; Li et al., 2021), based on the expectation that blue galaxies do not show significant IA. This work shows a persisting alignment of blue galaxies in most measurement cases.
@@ 5 Conclusions
    The distributions of the rir-i, $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$ properties of galaxies show variation in shape and redshift evolution across TNG300, EAGLE and Horizon-AGN. For rir-i, the bimodality of the distribution in TNG300, or unimodality in TNG300 and Horizon-AGN is the largest discrepancy. For $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$ it is the opposing redshift evolution between EAGLE and the other two simulations (Fig. 1-4).
@@
    The stellar mass distributions of the elliptical samples are fairly consistent across simulations, but vary between morphological selections: where selecting by $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$ include many low-mass galaxies ($M_{\star}\lesssim 10^{10}\mathrm{M}_{\odot}/h$) and rir-i and selecting by $\mathrm{BTR}$ includes mainly higher mass galaxies. (Fig. 5)
@@
The catalogues with the variables described in Section 2.4 and the galaxy shapes used in this work will be publicly available upon acceptance. The correlations can be measured using these catalogues and the public code MeasureIA.
@@ Acknowledgements
Appendix A Resolution and boxsize
@@
Figure 13: The quadrupoles of ellipticals (continuous lines) and disks (dashed) around all resolved galaxies in TNG100 at $z=0$ (left) and $z=1$ (right) for shapes calculated using the simple (top) or reduced (bottom) inertia tensor. The shape samples are defined by the values of the $|v/\sigma|$ (light blue), $\kappa_{rot}$ (dark blue), rir-i (orange) or $M_{\star}$ (green) of the galaxies. The results are similar to those of TNG300, albeit noisier.
@@ Appendix B $w_{g+}$ results
Appendix C Disks and ellipticals around galaxies: EAGLE and Horizon-AGN
Figure 15: The quadrupoles of disks (dashed) and ellipticals (continuous) around galaxies for various morphological definitions (colour), for simple shapes at $z=0$ (first column), simple shapes at $z=1$ (second column), reduced shapes at $z=0$ (third column), reduced shapes at $z=1$ (fourth column), in EAGLE (bottom row), Horizon-AGN (middle row), and TNG300 (top row). The shape samples are defined by the values of the $|v/\sigma|$ (light blue), $\kappa_{rot}$ (dark blue), rir-i (orange) or $M_{\star}$ (green) of the galaxies. The main trends persist across redshift and shape definition for each simulation.
@@ Appendix C Disks and ellipticals around galaxies: EAGLE and Horizon-AGN
Appendix D Bulge-to-total ratio

### v2

@@ 4 Discussion & Outlook
Our work goes beyond the existing literature by remeasuring the shapes and variables used in the morphology definitions in a consistent way between simulations. Although we aim to make a fair comparison between the different simulations, due to the variations in the distributions of the morphologically defining variables, as described in Section 3.1, finding a consistent way to define and compare the samples across simulations remains a challenge.
@@
While this work focusses on the direct comparison of the intrinsic alignment correlations between and within simulations, the origin of the differences between the distributions of the variables and their redshift evolution and between the correlations themselves remains unclear, as underlying stellar mass distributions of the samples, boxsize of the simulation, resolution, sub-grid physics models and galaxy and halo finder algorithms could all be playing a role. Currently, we are unable to distinguish between physical phenomena and simulation artifacts. Without a clear preference from observations or a better understanding of the effects of modelling choices in simulations on these statistics, the interpretation of the results remains limited. From theory, we do have some expectation on the amplitude of the intrinsic alignment correlation functions, which are given in the works of Camelio & Lombardi (2015) and Ghosh et al. (2024) who both estimate the quadrupolar distortion of the phase-space distribution of the stars in a galaxy, created by an external tidal field. Ghosh et al. (2024) distinguish between ellipticals and disks, and also estimate how the alignment amplitude of these galaxies depend on properties such as mass, size and Sérsic-index. While both studies find lower linear alignment amplitudes than measured in observations, they provide a theoretical, physical motivation for the alignment correlations found in simulations and observations. At higher order, Codis et al. (2015) attempts to determine the influence of tidal torques on the angular momentum alignment of disk-like galaxies from theory. Their findings suggest that below a certain mass threshold, angular momenta align parallel to filaments, and perpendicular above it. This could play a role in the interpretation of the sign of the reduced shape alignment signal, although it does not justify the differences found between hydrodynamical simulations.
@@
The $\mathrm{BTR}$ (Appendix D) and $|v/\sigma|$ splits in this work are similar to the work of Tenneti et al. (2016), where both variables are used as a morphological split to measure the ellipticity-direction correlations of disks around ellipticals in MassiveBlack-II (Khandai et al., 2015) and Illustris (Genel et al., 2014; Vogelsberger et al., 2014). These measurements are compared to those in Chisari et al. (2015), where $|v/\sigma|$ is used as a morphological proxy to measure the IA correlations of disks around ellipticals in Horizon-AGN. Tenneti et al. (2016) measure a positive (radial) disk around elliptical correlation in MassiveBlack-II and Illustris for selections based on both $\mathrm{BTR}$ and $|v/\sigma|$ and show that while the amplitude changes with disk definition, the sign is always positive.
@@
Although this work gives insight into under what circumstances a negative disk signal is measured, further sub-selections of the galaxy samples and larger statistics could help unveil this. Measuring the same signal in larger boxes, such as the Horizon Run 5 (Lee et al., 2021), COLIBRE (Schaye et al., 2025; Chaikin et al., 2025) and MilleniumTNG (Delgado et al., 2023), would also be very useful to continue this research. The effects of boxsize and resolution could also be explored further within these projects, as there will be multiple boxsizes and resolutions available that meet the minimum requirements to measure statistically significant correlations. To further this research area, future hydrodynamical simulations could focus on obtaining colours that are directly comparable to observations, making it easier to mimic the selections of observational samples. Furthermore, it would be beneficial if galaxy kinematics are more realistic and understood.
Another area to expand this research to would be the observational realm. Currently, little is known about how alignments respond to the kinematic properties of galaxies, as the measurement of $|v/\sigma|$ are very difficult to obtain observationally and only available for small samples. Expanding works like Pulsoni et al. (2020) that compare kinematic properties of galaxies between hydrodynamical simulations and observations for small samples, will provide valuable insight on the accuracy of the models we currently employ.
This work highlights the importance of consistency when comparing between different disk and elliptical alignments under varying circumstances. Although hydrodynamical simulations agree on many features of the alignment signal, details like sample selection choices or measurement methodologies can make a substantive difference in outcome. Furthermore, this works calls for caution in weak lensing studies that restrict to blue galaxies to mitigate IA effects (e.g. Li et al., 2021; McCullough et al., 2024), based on the expectation that blue galaxies do not show significant IA. This work shows a persisting alignment of blue galaxies in most measurement cases.
@@ 5 Conclusions
    The distributions of the rir-i, $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$ properties of galaxies show variation in shape and redshift evolution across TNG300, EAGLE and Horizon-AGN. For rir-i, the bimodality of the distribution in TNG300, or unimodality in EAGLE and Horizon-AGN is the largest discrepancy. For $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$ it is the opposing redshift evolution between EAGLE and the other two simulations (Fig. 1-4).
@@
    The stellar mass distributions of the elliptical samples are fairly consistent across simulations, but vary between morphological selections: where selecting by $|v/\sigma|$ or $\kappa_{\mathrm{rot}}$ include many low-mass galaxies ($M_{\star}\lesssim 10^{10}\mathrm{M}_{\odot}/h$) and selecting by rir-i or $\mathrm{BTR}$ includes mainly higher mass galaxies. (Fig. 5)
@@
The catalogues with the variables described in Section 2.4 and the galaxy shapes used in this work are publicly available for TNG300888https://doi.org/10.5281/zenodo.18740718, EAGLE [^9] https://doi.org/10.5281/zenodo.18740922 and Horizon-AGN [^10] https://doi.org/10.5281/zenodo.18740433. The correlations can be measured using these catalogues and the public code MeasureIA  [^11] https://github.com/MarloesvL/measure_IA121212doi:10.5281/zenodo.17252215.
@@ Acknowledgements
Appendix A Appendix A: Resolution and boxsize
@@
Figure 13: The quadrupoles of ellipticals (continuous lines) and disks (dashed) around all resolved galaxies in TNG100 at $z=0$ (left) and $z=1$ (right) for shapes calculated using the simple (top) or reduced (bottom) inertia tensor. The shape samples are defined by the values of the $|v/\sigma|$ (light blue), $\kappa_{rot}$ (dark blue), rir-i (orange) or $M_{\star}$ (green) of the galaxies. The results are similar to those of TNG300, given in Fig. 7, albeit noisier.
@@ Appendix B Appendix B: $w_{g+}$ results
Appendix C Appendix C: Disks and ellipticals around galaxies: EAGLE and Horizon-AGN
Figure 15: The quadrupoles of disks (dashed) and ellipticals (continuous) around galaxies for various morphological definitions (colour), for simple shapes at $z=0$ (first column), simple shapes at $z=1$ (second column), reduced shapes at $z=0$ (third column), reduced shapes at $z=1$ (fourth column), in EAGLE (bottom row), Horizon-AGN (middle row), and TNG300 (top row, same as Fig. 7). The shape samples are defined by the values of the $|v/\sigma|$ (light blue), $\kappa_{rot}$ (dark blue), rir-i (orange) or $M_{\star}$ (green) of the galaxies. The scale dependence and ordering of the amplitudes of the different samples persist across redshift and shape definition for each simulation.
@@ Appendix C Appendix C: Disks and ellipticals around galaxies: EAGLE and Horizon-AGN
Appendix D Appendix D: Bulge-to-total ratio